OpenStudy (happykiddo):

Find the center, foci, and vertices of the hyperbola. Use a graphing utility to graph the hyperbola and its asymptotes. 16y2 − x2 + 2x + 128y + 175 = 0

OpenStudy (anonymous):

\[16y^2 − x^2 + 2x + 128y + 175 = 0\] subtract \(175\) and get \[16y^2+128y − x^2 + 2x =-1750\] then time to complete the square

OpenStudy (anonymous):

damn

OpenStudy (happykiddo):

I got the center, I need help finding the foci and vertices.

OpenStudy (anonymous):

\[16y^2+128y-x^2+2x=-175\]

OpenStudy (anonymous):

did you get in in the form \[\frac{(y-k)^2}{a^2}-\frac{(x-h)^2}{b^2}=1\]?

OpenStudy (happykiddo):

Yeah I know the formula but I get sqrt 5=a and 2sqrt5=b. And it says its wrong